HUBBLE SPACE TELESCOPE DAILY REPORT #1115 PERIOD COVERED: 0700 EST Daily Status Report as of 102/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: The spacecraft entered software safemode at 101/0559Z. All observations after that time did not execute, and were mark as "failed". 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 0 Successful : 0 Scheduled Reacquisitions : 0 Successful : 0 2.2 FHST Updates: Scheduled :0 Successful :0 2.3 Significant HSTARs: None 2.4 Operations Notes: The observatory remains in software sunpoint safemode. The anomaly investigation team continues to meet. The Gimbal Electronic Assemblies and the FHSTs were powered on immediately following communications recovery yesterday morning. The DCE was powered on at 1300Z. The PSEA was reconfigured for SADE 2 operations at 1405Z. FHST maps and tape recorder dumps are currently being managed by real-time commands. The tentative recovery schedule specifies that the health and safety SMS will begin this evening at 103/0000Z, and the first SMS from the STScI will begin execution tomorrow afternoon at 103/1900Z. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: The first SMS from the STScI will begin execution tomorrow afternoon at 103/1900Z. 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Recovery of the spacecraft from software sunpoint safemode, and the resumption of normal science observations. /DMH -END-